# Infrared finite scattering theory: Scattering states and representations of the BMS group

### v1

@@ 2 Orbit spaces of memory under Lorentz transformations
where the functions $\varpi_{\Lambda}$ and $\phi_{\Lambda}$ are given explicitly in Appendix A. The power of $\varpi_{\Lambda}$ and $e^{i\phi_{\Lambda}}$ appearing in the Lorentz transformation eq. 3.25 refer to the fact that the complex function $\Delta$ has conformal weight $w=-1$ and spin weight $s=-2$ (see appendix A for the definition of these weights).
@@
Another simple case of physical interest is when the memory is non-vanishing but has no symmetries. This is the generic case that is expected to occur for any non-trivial scattering, e.g., if the in-state consists of an $n$-graviton state with vanishing memory then the out-state will generically have a memory that has no symmetries. A more precise argument to this effect will be given in a forthcoming paper [42]. In this case, the little group is the subgroup $\mathbb{Z}_{2}$ (consisting of the identity matrix and its negative) which has trivial action on $\mathbb{S}^{2}$. The orbit space is the identity component of the Lorentz group ${\tt SO}(3,1)$ which is homeomorphic to $\mathbb{R}^{3}\times\mathbb{R}\mathbb{P}^{3}$ where $\mathbb{R}\mathbb{P}^{3}$ is the sphere $\mathbb{S}^{3}$ with the antipodal points identified. Coordinates on this orbit space can be obtained directly from suitable coordinates on the Lorentz group (see eq. A.28). These are Cartesian coordinates $(t,u,v)$ on $\mathbb{R}^{3}$ and standard spherical coordinates $(\theta,\phi,\psi)$ on $\mathbb{S}^{3}$. The Lorentz invariant measure on this orbit space is
@@ Acknowledgements.
G.S. is supported by the Princeton Gravity Initiative at Princeton University. This work was supported in part the NSF grant PHY-2107939 to the University of California, Santa Barbara.
@@ Appendix A Little groups and orbits spaces of conformal-spin weighted functions on the $2$-sphere
where we have used the fact that, from Appendix A, we have $q_{AB}(z,\overline{z})=\varpi_{\Lambda}^{-2}q_{AB}(\Lambda z,\overline{\Lambda z})$ and $m_{A}(z,\overline{z})=\varpi_{\Lambda}^{-1}e^{-i\phi_{\Lambda}}m_{A}(\Lambda z,\overline{\Lambda z})$ Henceforth we do not write the explicit dependence of the function on the metric and basis for notational convenience, and denote the conformal-spin weights of a function by $(w,s)$.
@@
The inversion of a function $\varphi(z,\overline{z})$ of weight $(w,s)$ is then given by (using Appendices A and A.11)
@@
Using eqs. A.2, A and A.11, the transformation of $\hat{\varphi}(z,\overline{z})$ under the action of $\Lambda\in{\tt SL}(2,\mathbb{C})$ is given by
@@ 1 Construction of the little groups and orbit spaces
Figure 1: Relations between the connected subgroups in the notation of [47]. The arrows indicate inclusion (up to conjugation) as a subgroup, e.g. $\Gamma\rightarrow P$ indicates that $P$ contains a subgroup conjugate to $\Gamma$. The center $\mathbb{Z}_{2}$ and other disconnected subgroups have not been included. See also footnote 18.
@@ 3 Non-connected little groups
We briefly remark on the possible non-connected subgroups of ${\tt SL}(2,\mathbb{C})$ that arise as possible little groups. The case of the center $\mathbb{Z}_{2}$ has already been considered since it acts trivially on $\mathbb{S}^{2}$. Following, McCarthy [44] it can be shown that the only other non-connected groups which arise as possible little groups are the lifts to ${\tt SU}(2)$ of non-connected subgroups of the rotation group ${\tt SO}(3)$. These are considered below.

### v2

@@ 2 Orbit spaces of memory under Lorentz transformations
where the functions $\varpi_{\Lambda}$ and $\phi_{\Lambda}$ are given explicitly in eq. A.7. The power of $\varpi_{\Lambda}$ and $e^{i\phi_{\Lambda}}$ appearing in the Lorentz transformation eq. 3.25 refer to the fact that the complex function $\Delta$ has conformal weight $w=-1$ and spin weight $s=-2$ (see appendix A for the definition of these weights).
@@
Another simple case of physical interest is when the memory is non-vanishing but has no symmetries. This is the generic case that is expected to occur for any non-trivial scattering, e.g., if the in-state consists of an $n$-graviton state with vanishing memory then the out-state will generically have a memory that has no symmetries. A more precise argument to this effect is given in [44]. In this case, the little group is the subgroup $\mathbb{Z}_{2}$ (consisting of the identity matrix and its negative) which has trivial action on $\mathbb{S}^{2}$. The orbit space is the identity component of the Lorentz group ${\tt SO}(3,1)$ which is homeomorphic to $\mathbb{R}^{3}\times\mathbb{R}\mathbb{P}^{3}$ where $\mathbb{R}\mathbb{P}^{3}$ is the sphere $\mathbb{S}^{3}$ with the antipodal points identified. Coordinates on this orbit space can be obtained directly from suitable coordinates on the Lorentz group (see eq. A.28). These are Cartesian coordinates $(t,u,v)$ on $\mathbb{R}^{3}$ and standard spherical coordinates $(\theta,\phi,\psi)$ on $\mathbb{S}^{3}$. The Lorentz invariant measure on this orbit space is
@@ Acknowledgements.
G.S. is supported by the Princeton Gravity Initiative at Princeton University. This work was supported in part by the NSF grant PHY-2107939 to the University of California, Santa Barbara.
@@ Appendix A Little groups and orbits spaces of conformal-spin weighted functions on the $2$-sphere
where we have used the fact that, from eq. A.6, we have $q_{AB}(z,\overline{z})=\varpi_{\Lambda}^{-2}q_{AB}(\Lambda z,\overline{\Lambda z})$ and $m_{A}(z,\overline{z})=\varpi_{\Lambda}^{-1}e^{-i\phi_{\Lambda}}m_{A}(\Lambda z,\overline{\Lambda z})$ Henceforth we do not write the explicit dependence of the function on the metric and basis for notational convenience, and denote the conformal-spin weights of a function by $(w,s)$.
@@
The inversion of a function $\varphi(z,\overline{z})$ of weight $(w,s)$ is then given by (using eqs. A.7 and A.11)
@@
Using eqs. A.2, A.7 and A.11, the transformation of $\hat{\varphi}(z,\overline{z})$ under the action of $\Lambda\in{\tt SL}(2,\mathbb{C})$ is given by
@@ 1 Construction of the little groups and orbit spaces
Figure 1: Relations between the connected subgroups in the notation of [49]. The arrows indicate inclusion (up to conjugation) as a subgroup, e.g. $\Gamma\rightarrow{\tt P}$ indicates that ${\tt P}$ contains a subgroup conjugate to $\Gamma$. The center $\mathbb{Z}_{2}$ and other disconnected subgroups have not been included. See also footnote 19.
@@ 3 Non-connected little groups
We briefly remark on the non-connected subgroups of ${\tt SL}(2,\mathbb{C})$ that arise as possible little groups. The case of the center $\mathbb{Z}_{2}$ has already been considered since it acts trivially on $\mathbb{S}^{2}$. Following, McCarthy [46] it can be shown that the only other non-connected groups which arise as possible little groups are the lifts to ${\tt SU}(2)$ of non-connected subgroups of the rotation group ${\tt SO}(3)$. These are considered below.